Determine the values of P x and Ex for each of the following signals: (a) x 1 (t) = e- 21 u(t) 


1
Expert's answer
2021-05-21T04:04:01-0400

"x_1(t)=e^{-2t}u(t)"

"t=\infin \implies x_1(t)=0 \implies" It is energy signal 

Therefore, Power , Px=0

Energy , "Ex=\lim\limits_{t\to \infin} \int_0^t(e^{-2t})^2dt=\lim\limits_{t\to \infin} \frac{e^{-ut}}{-u}|_0^t=\frac{1}{4}J"